acetylen,c2h2 is used in weldng torches. It releases a lot of energy when burned in oxygen. One gram of acetylene releases 48.2kJ. When 0.75g of acetylene is burned in a bomb calorimeter (heat capacity=1.117kJ/°C), the final temperaturre of the bomb is 54.5°C. What is the initial temperature of the calorimeter? (Pls provide solutions) Thank you
1
Expert's answer
2019-06-28T02:46:27-0400

Solution.

If during the burning of 1 gram of acetylene 48.2 kJ is released, then during the burning of 0.75 grams it is released:


"Q = \frac{0.75 \times 48.2}{1}"

Q = 36.15 kJ


"Q = q \times \Delta t"

q = 1.117 kJ/°C


"\Delta t = \frac{Q}{q}""\Delta t = t2-t1""\Delta t = \frac{36.15}{1.117}"

"\Delta t = 32.36"

"t1 = t2 - \Delta t"

t1 = 22.14 °C

Answer:

t1 = 22.14 °C